OpenStudy (anonymous):

Evaluate the expression without using a calculator. log101000

OpenStudy (anonymous):

\[\log _{10}1000\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

10^3= 1000

OpenStudy (anonymous):

Oh, cool. Thanks

OpenStudy (anonymous):

log 10 1000 = x 10^x=1000 x=3

OpenStudy (cathyangs):

so for \[\log_{x} y\]=z x^z=y

OpenStudy (anonymous):

yes